# Safeguarding the Truth of High-Value Price Oracle Task: A Dynamically Adjusted Truth Discovery Method

### 04 Feb 2024 v1

@@
Abstract
@@
Index Terms:
Blockchain, Oracle, DeFi, Truth Discovery.
I Introduction
@@
II Preliminary and Existing Problems
@@ III The proposed scheme
III-A Overview
Specifically, considering the two key problems that exist at present, we design a new truth discovery method to protect the truth of the high-value price oracle. Compared with the traditional TD algorithm, our method is mainly optimised from the two stages of truth aggregation and credibility evaluation to improve the resistance of the TD algorithm to high-value attacks, which are introduced in detail in \$III-B and \$III-C.
@@ III-A Overview
    Return data: The oracle contract calls the callback function $F$, and returns the final truth of the aggregation $\hat{x}^{I}$ to the user contract that invokes the interface.
@@
III-B Truth Aggregation
Input: The aggregation weight set of $S$, the data set obtained from $S$, and task value in round $I$: $\{x_{s}^{I}\},\{w_{s}^{I}\}$, $M^{I}$
@@ III-B Truth Aggregation
1 $\hat{x}^{I}=\frac{\sum_{s\in S}{\mathbb{1}_{s}^{I}\cdot w_{s}\cdot x_{s}^{I}}}{\sum_{s\in S}{\mathbb{1}_{s}^{I}\cdot w_{s}}}$;
@@ 14       $r_{s}^{I+1}=Sigmod(\varrho_{s}\frac{\sum_{i=1}^{I}M^{i}d(x_{s}^{I},\hat{x}^{I})}{M^{I}})$ ;
To solve problem 1 (\$II-C1), we use the idea of the temporal difference algorithm to increase the consideration of the future in the original truth aggregation stage, so that the estimation of the truth is more accurate. Because the source is credible in the past, it does not mean that the present and the future are credible. In the traditional TD method, if the source performs poorly in this task, its loss will only be reflected after this credibility update, but the truth aggregation stage of this round will also be influenced by its previous high credibility.
@@
We distinguish the two-stage TD in the Table. I. Both \$III-B and \$III-C are based on the first TD, and the process of the second TD is similar, so we will not repeat it too much.
TABLE I: The difference between two-stage TD.
@@
III-C Credibility Update
To solve question 2 (\$II-C2), we added the consideration of task value, participation rate, etc. to the process of credibility evaluation. In the traditional TD method, if the source is malicious, the credibility of the loss is only related to the degree of data deviation, and has nothing to do with the actual economic loss of the task. A malicious user can easily get the value of the task from the blockchain, and can cause a lot of economic loss by doing bad things to high-value tasks. But they can restore their credibility by behaving honestly on a large number of low-value tasks, which is obviously unreasonable.
@@ III-C Credibility Update
$r_{s}^{I+1}=Sigmod(\varrho_{s}\frac{\sum_{i=1}^{I}M^{i}d(x_{s}^{I},\hat{x}^{I})}{M^{I}})\ \forall s\in S$ (12)
@@
III-D Examples
TABLE II: Some parameters involved in the process of truth discovery from $S$ by $O_{1}$.
@@ III-D Examples
Parameter Setting The discount factor is set to $\gamma=0.5$, with the current task value being $M^{I}=8$. The CPEC is calculated as $\vec{\mathcal{C}_{s}^{I}}=2.5$, and the cumulative participation rate is $\varrho_{s}=1.0$. Refer to Table. II for additional details.
As indicated in Table. II, we assume that the credibility $r_{s}^{I}$ of source $s$ in the initial task is $\{0.8,0.8,0.8,0.95,0.95\}$, with $s_{4}$ and $s_{5}$ being highly trusted sources. If this moment, the sources $s_{4}$ and $s_{5}$ suddenly defect and deliberately submit lower price data $x_{s}^{I}$ for profit, then according to Eq. (5), the truth $\hat{x}^{I}_{1}$ aggregated by $O_{1}$ using the conventional TD scheme is 0.757.
@@
III-E Analyses
Under the assumption specified in \$III-A, we demonstrate the data credibility of the proposed scheme under high-value attacks and Freeloading problems.
@@ III-E Analyses
Proof Sketch: When a source $s\in S$ succumbs to the temptation of high-value tasks and chooses to behave maliciously, its returned data deviates from the truth. In the truth aggregation stage, the aggregation weight $w_{s}$ not only considers the historical credibility $r_{s}^{I}$ of $s$ but also estimates its credibility in the next round $\tilde{r}_{s}^{I+1}$ based on its current performance. Even if $s$ has maintained high credibility in the past, due to dishonest behavior in the current task, according to Eq. (3), a very low $\tilde{r}_{s}^{I+1}$ will be assigned, resulting in a low final $w_{s}$. This effectively weakens the impact of $s$ on the current truth aggregation and subsequent tasks, as shown in III-D. It is essential to note that the accuracy of the TD algorithm is based on the assumption that the majority of sources are honest [23]. This ensures that in the truth aggregation stage of the first TD, there is always a probability of more than half of being close to the truth, and in the second TD, there is always a probability of more than $1-f_{o}$ of being close to the truth.
@@
IV Experiment and Result Analysis
(a) Data deviation under 10% attack
(b) Data deviation under 30% attack
(c) Potential economic losses under 10% attack
@@ IV Experiment and Result Analysis
Experiment Setting: In the established oracle network, there are 20 data sources and 20 oracle nodes. The proportions of malicious data sources and oracle nodes are denoted as $\alpha$ and $\beta$, respectively. The discount factor is set to $\gamma=0.5$. The task of each oracle is to acquire random price data with values ranging from 0 to 100. High-value tasks have values between 100 and 10,000, while low-value tasks range from 1 to 100. Malicious sources maintain honesty in low-value tasks, while randomly adjusting price data in high-value tasks by (0-50%) to seek profits. The baseline refers to the prototype implementation of DTDO [23].
IV-A Data Deviation
@@ IV-B Aggregation Weight and Credibility
Fig. 4 presents the aggregation weights of different schemes in each round of the truth aggregation stage along with their variations. The aggregation weight ratio in the graph indicates the ratio of the aggregation weight of an honest node to that of a malicious node. A higher ratio signifies a more accurate final aggregation result. Notably, in high-value tasks (as identified in the marker), the proposed scheme consistently maintains a higher weight ratio compared to the baseline. Furthermore, it is observed that the baseline consistently exhibits a slow reactive response to high-value attacks, potentially leading to misjudgments, aligning with our analysis in question 1 (\$II-C1). Lastly, under sustained malicious activity, such as around the 90th round in the figure, the proportion of the baseline rises rapidly, indicating an increased identification of malicious nodes. However, after 1-2 rounds of honest behavior, the proportion rapidly decreases, illustrating a swift forgetting of historical malicious behavior, in line with our analysis in question 2 (\$II-C2).
@@
IV-C Robustness
To assess the robustness of the proposed scheme, we conducted tests on data deviation under various conditions, including different schemes, malicious ratios $\beta$, malicious scenarios (modification range) $\omega$, and the first TD process. The results are presented in Fig. 6.
(a) Data deviation under different $\beta$
(b) Data deviation under different $\omega$
(c) Data deviation during the first TD process
Figure 6: The robustness of the proposed solution. Data deviation in (a) different malicious ratios $\beta$, (b) different malicious situations (modification scope) $\omega$, and (c) the first TD process.
Fig. 6a displays the data deviation of different schemes when confronted with high-value attacks at varying proportions of malicious nodes $\beta=$ [10%, 20%, 30%]. Notably, the proposed scheme consistently sustains a lower data deviation than the baseline and remains closer to the truth across different proportions of malicious nodes.
Xian et al. [25] suggested that malicious nodes might seek benefits by adjusting the scope of data tampering. Therefore, we conducted an analysis of data deviation when the range of malicious node data tampering, denoted as $\omega$, is (0-50%) and (0-200%) in Fig. 6b. The results reveal that even when malicious nodes employ different data tampering strategies, the proposed scheme consistently exhibits lower data deviation than the baseline.
@@ IV-C Robustness
Finally, to analyze data consistency with different values of $\gamma$ in the proposed scheme, we set $\gamma$ to [0.2, 0.4, 0.6, 0.8] and conduct 10 tasks. Experimental results indicate that when 10% of tasks are attacked by high-value attacks, data deviation increases with the rise of $\gamma$. This aligns with our assumption that increasing attention to the future enhances the node’s ability to respond to sudden attacks. However, when attacks are more frequent, such as when there are high-value attacks in 80% of tasks, proper consideration of historical credibility can help strengthen the identification of malicious nodes.
@@
V Conclusion
This paper introduces a dynamically adjusted truth discovery method aimed at safeguarding the truth of high-value price oracle tasks. To enhance resistance against high-value attacks in the truth aggregation stage, we incorporate the concept of temporal difference algorithms, increasing consideration of future reliability during the aggregation stage to minimize truth aggregation deviation. Subsequently, in the credibility update stage, we dynamically adjust credibility based on factors such as task value and cumulative potential economic contribution of the sources, improving source evaluation accuracy. Experimental results demonstrate that in the presence of high-value attacks, the proposed scheme exhibits lower data deviation compared to the baseline scheme and reduces potential economic losses.
@@ V Conclusion
Acknowledgments
The research was supported in part by the Guangxi Science and Technology Major Project (No. AA22068070), the National Natural Science Foundation of China (Nos. 62166004,U21A20474), the Basic Ability Enhancement Program for Young and Middle-aged Teachers of Guangxi (No.2022KY0057 ), the Key Lab of Education Blockchain and Intelligent Technology, the Center for Applied Mathematics of Guangxi, the Guangxi ”Bagui Scholar” Teams for Innovation and Research Project, the Guangxi Talent Highland Project of Big Data Intelligence and Application, the Guangxi Collaborative Center of Multisource Information Integration and Intelligent Processing.

### 22 Feb 2024 v2

@@
[…]
Safeguarding the Truth of High-Value Price Oracle Task: A Dynamically Adjusted Truth Discovery Method
(2015; 2015)
keywords:
Blockchain, Oracle, DeFi, Truth Discovery.
{summary}
@@
1 Introduction
@@
The rest of this paper is organized as follows. Section 2 introduces the oracle, truth discovery methods, and current problems. Section 3 introduces the proposed scheme and provides an example and analysis. Section 4 presents the experimental results and analysis. Finally, Section 5 concludes this paper and outlines ideas for future work.
2 Preliminary and Existing Problems
2.1 Oracle
@@ 3 The proposed scheme
3.1 Overview
Specifically, considering the two key problems that exist at present, we design a new truth discovery method to protect the truth of the high-value price oracle. Compared with the traditional TD algorithm, our method is mainly optimised from the two stages of truth aggregation and credibility evaluation to improve the resistance of the TD algorithm to high-value attacks, which are introduced in detail in \$3.2 and \$3.3.
@@ 3.1 Overview
    Data callback: The oracle contract calls the callback function $F$, and returns the final truth of the aggregation $\hat{x}^{I}$ to the user contract that invokes the interface.
@@
3.2 Truth Aggregation
Input: The aggregation weight set of $S$, the data set obtained from $S$, and task value in round $I$: $\{w_{s}^{I}\}$, $\{x_{s}^{I}\}$, $M^{I}$
@@ 3.2 Truth Aggregation
1 $\hat{x}^{I}_{i}=\frac{\sum_{s\in S}{\mathbb{1}_{s}^{I}\cdot w_{s}\cdot x_{s}^{I}}}{\sum_{s\in S}{\mathbb{1}_{s}^{I}\cdot w_{s}}}$;
@@ 14       $r_{s}^{I+1}=Sigmod(\varrho_{s}\frac{\sum_{i=1}^{I}M^{i}d(x_{s}^{I},\hat{x}^{I})}{M^{I}})$ ;
To solve question 1 (\$2.3.1), we use the idea of the temporal difference algorithm to increase the consideration of the future in the original truth aggregation stage, so that the estimation of the truth is more accurate. Because the source is credible in the past, it does not mean that the present and the future are credible. In the traditional TD method, if the source performs poorly in this task, its loss will only be reflected after this credibility update, but the truth aggregation stage of this round will also be influenced by its previous high credibility.
@@
We distinguish the two-stage TD in the Table. 1. Both \$3.2 and \$3.3 are based on the first TD, and the process of the second TD is similar, so we will not repeat it too much.
Table 1: The difference between two-stage TD.
@@
3.3 Credibility Update
To solve question 2 (\$2.3.2), we added the consideration of task value, participation rate, etc. to the process of credibility evaluation. In the traditional TD method, if the source is malicious, the credibility of the loss is only related to the degree of data deviation, and has nothing to do with the actual economic loss of the task. A malicious user can easily get the value of the task from the blockchain, and can cause a lot of economic loss by doing bad things to high-value tasks. But they can restore their credibility by behaving honestly on a large number of low-value tasks, which is obviously unreasonable.
@@ 3.3 Credibility Update
$r_{s}^{I+1}=Sigmod(\varrho_{s}\cdot\frac{\vec{\mathcal{C}_{s}^{I}}}{M^{I}})\ \forall s\in S$ (12)
@@
3.4 Examples
Parameter Setting: The discount factor is set to $\gamma=0.5$, with the current task value being $M^{I}=8$. The CPEC is calculated as $\vec{\mathcal{C}_{s}^{I}}=2.5$, and the cumulative participation rate is $\varrho_{s}=1.0$. Refer to Table. 2 for additional details.
Table 2: Some parameters involved in the process of truth discovery from $S$ by $O_{1}$.
@@ 3.4 Examples
As indicated in Table. 2, we assume that the credibility $r_{s}^{I}$ of source $s$ in the initial task is $\{0.8,0.8,0.8,0.95,0.95\}$, with $s_{4}$ and $s_{5}$ being highly trusted sources. If this moment, the sources $s_{4}$ and $s_{5}$ suddenly defect and deliberately submit lower price data $x_{s}^{I}$ for profit, then according to Eq. (5), the truth $\hat{x}^{I}_{1}$ aggregated by $O_{1}$ using the conventional TD scheme is 0.757.
@@
3.5 Analyses
Under the assumption specified in \$3.1, we demonstrate the data credibility of the proposed scheme under high-value attacks and Freeloading problems.
@@ 3.5 Analyses
Proof Sketch: When a source $s\in S$ succumbs to the temptation of high-value tasks and chooses to behave maliciously, its returned data deviates from the truth. In the truth aggregation stage, the aggregation weight $w_{s}$ not only considers the historical credibility $r_{s}^{I}$ of $s$ but also estimates its credibility in the next round $\tilde{r}_{s}^{I+1}$ based on its current performance. Even if $s$ has maintained high credibility in the past, due to dishonest behavior in the current task, according to Eq. (3), a very low $\tilde{r}_{s}^{I+1}$ will be assigned, resulting in a low final $w_{s}$. This effectively weakens the impact of $s$ on the current truth aggregation and subsequent tasks, as shown in 3.4. It is essential to note that the accuracy of the TD algorithm is based on the assumption that the majority of sources are honest [23]. This ensures that in the truth aggregation stage of the first TD, there is always a probability of more than half of being close to the truth, and in the second TD, there is always a probability of more than $1-f_{o}$ of being close to the truth.
@@
(a) Data deviation when $\tau$=0.1
(b) Data deviation when $\tau$=0.3
(c) Potential economic losses when $\tau$=0.1
@@ 4 Experiment and Result Analysis
Experiment Setting: In the established oracle network, there are 20 data sources and 20 oracle nodes. The proportions of malicious data sources and oracle nodes are denoted as $\alpha$ and $\beta$, respectively. The discount factor is set to $\gamma=0.5$. The task of each oracle is to acquire random price data with values ranging from 0 to 100. High-value tasks have values between 100 and 10,000, while low-value tasks range from 1 to 100. Malicious sources maintain honesty in low-value tasks, while randomly adjusting price data in high-value tasks by (0-50%) to seek profits. The baseline refers to the prototype implementation of DTDO [23]. The experimental parameters are set in Table.3.
Table 3: Experimental parameter settings.
notation meaning value
$\gamma$ discount factor 0.5
$\alpha$ proportion of malicious data sources 0.4
$\beta$ proportion of malicious oracle nodes 0.3
$\omega$ range of malicious node data tampering (0-50%)
$\tau$ proportion of high-value attacks 0.1
4.1 Data Deviation
@@
(a) Data deviation with different $\beta$
(b) Data deviation with different $\omega$
(c) Data deviation in the first TD
Figure 6: The robustness of the proposed solution. Data deviation in (a) different proportion of malicious nodes $\beta$, (b) different range of malicious node data tampering $\omega$, and (c) the first TD.
4.2 Aggregation Weight and Credibility
Fig. 4 presents the aggregation weights of different schemes in each round of the truth aggregation stage along with their variations. The aggregation weight ratio in the graph indicates the ratio of the aggregation weight of an honest node to that of a malicious node. A higher ratio signifies a more accurate final aggregation result. Notably, in high-value tasks (as identified in the marker), the proposed scheme consistently maintains a higher weight ratio compared to the baseline. Furthermore, it is observed that the baseline consistently exhibits a slow reactive response to high-value attacks, potentially leading to misjudgments, aligning with our analysis in question 1 (\$2.3.1). Lastly, under sustained malicious activity, such as around the 90th round in the figure, the proportion of the baseline rises rapidly, indicating an increased identification of malicious nodes. However, after 1-2 rounds of honest behavior, the proportion rapidly decreases, illustrating a swift forgetting of historical malicious behavior, in line with our analysis in question 2 (\$2.3.2).
@@ 4.2 Aggregation Weight and Credibility
4.3 Robustness
To assess the robustness of the proposed scheme, we conducted tests on data deviation under various conditions, including different schemes, malicious nodes ratios $\beta$, malicious strategy (modification range) $\omega$, and the first TD process. The results are presented in Fig. 6.
Fig. 6a displays the total data deviation of different schemes when confronted with high-value attacks at varying proportions of malicious nodes $\beta=$ [10%, 20%, 30%]. Notably, the proposed scheme consistently sustains a lower data deviation than the baseline and remains closer to the truth across different proportions of malicious nodes.
Xian et al. [25] suggested that malicious nodes might seek benefits by adjusting the scope of data tampering. Therefore, we conducted an analysis of total data deviation when the range of malicious node data tampering, denoted as $\omega$, is (0-50%), (0-100%) and (0-200%) in Fig. 6b. The results reveal that even when malicious nodes employ different data tampering strategies, the proposed scheme consistently exhibits lower data deviation than the baseline.
@@ 4.3 Robustness
Finally, to analyze data consistency with different values of $\gamma$ in the proposed scheme, we set $\gamma$ to [0.2, 0.4, 0.6, 0.8] and calculate the total data deviation for 10 tasks. Experimental results indicate that when 10% of tasks are attacked by high-value attacks, data deviation increases with the rise of $\gamma$. This aligns with our assumption that increasing attention to the future enhances the node’s ability to respond to sudden attacks. However, when attacks are more frequent, such as when there are high-value attacks in 80% of tasks, proper consideration of historical credibility can help strengthen the identification of malicious nodes. Finally, no matter how the environment changes, the proposed scheme can always provide better data than baseline.
5 Conclusion
This paper introduces a dynamically adjusted truth discovery method aimed at safeguarding the truth of high-value price oracle tasks. To enhance resistance against high-value attacks in the truth aggregation stage, we incorporate the concept of temporal difference algorithms, increasing consideration of future credibility during the aggregation stage to minimize truth aggregation deviation. Subsequently, in the credibility update stage, we dynamically adjust credibility based on factors such as task value and cumulative potential economic contribution of the sources, improving source evaluation accuracy. Experimental results demonstrate that in the presence of high-value attacks, the proposed scheme exhibits lower data deviation compared to the baseline scheme and reduces potential economic losses.
@@ 5 Conclusion
Acknowledgements
The research was supported in part by the Guangxi Science and Technology Major Project (No. AA22068070), the National Natural Science Foundation of China (Nos. 62166004,U21A20474), the Basic Ability Enhancement Program for Young and Middle-aged Teachers of Guangxi (No.2022KY0057, 2023KY0062), the Key Lab of Education Blockchain and Intelligent Technology, the Center for Applied Mathematics of Guangxi, the Guangxi ”Bagui Scholar” Teams for Innovation and Research Project, the Guangxi Talent Highland Project of Big Data Intelligence and Application, the Guangxi Collaborative Center of Multisource Information Integration and Intelligent Processing.